A certain weak acid (pKa = 5.2) is used as a food preservative. Determine the pH of a 0.54 M solution.


1
Expert's answer
2021-09-03T06:04:08-0400

Let the weak acid be HA

HA "\to" H+ + A-

at t =0 0.54_________-_____-

at t = teq 0.54-x _____x___x


Ka = "\frac{[H^+][A^-]}{[HA]}"


10-5.2 = "\frac{[x][x]}{[0.54-x]}"

0.54-x "\mapsto" 0.54 since x is very less


0.54×10-5.2 = x2

10-0.26 × 10-5.2 = x2

10-5.46 = x2

1.86 × 10-3 = x

[H+] = 1.86 ×10-3

[H+] = 10-2.73

pH = 2.73